OpenStudy (anonymous):

Perform the operations on complex numbers: (5+2i) / (3+6i)

hartnn (hartnn):

Multiply numerator and denominator with the complex conjugate of 3+6i

hartnn (hartnn):

you know how to do that?

OpenStudy (anonymous):

Yes, so multiply by (3-6i)?

hartnn (hartnn):

correct

hartnn (hartnn):

for the denominator, you can use \(\Large (a+bi)(a-bi) = a^2+b^2\)

hartnn (hartnn):

for the numerator, multiply the 2 factors out \((5+2i)(3-6i) = ...\)

OpenStudy (anonymous):

thank you! So the numerator is 15-12i?

OpenStudy (anonymous):

@hartnn, you keep beating me to all of the questions! Lol

hartnn (hartnn):

not actually... \(\Large (5+2i)(3-6i) = 5\times 3 + 5\times (-6i)+(2i)\times 3+(2i)\times (-6i)\)

hartnn (hartnn):

\(\large (5+2i)(3-6i) = 5\times 3 + 5\times (-6i)+(2i)\times 3+(2i)\times (-6i)\)

hartnn (hartnn):

@Jamal5337 If you want, you can continue helping him/her :)

OpenStudy (anonymous):

It's okay, I'm helping JonnyChewy now!

OpenStudy (anonymous):

Thanks hartnn!

hartnn (hartnn):

did you get the required answer? :)

hartnn (hartnn):

oh and bdw \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\)

OpenStudy (anonymous):

Another approach: Convert the given complex numbers to polar form. \[\large\frac{z}{w}=\frac{re^{i\alpha}}{se^{i\beta}}=\frac{r}{s}e^{i(\alpha-\beta)}\]